Sharks' teeth are shaped like triangles. The height of a tooth of a certain shark is 1 cm longer than the base. The area is 21 cm^2. Find the height and base.
Answer by nerdybill(7384)

Sharks' teeth are shaped like triangles. The height of a tooth of a certain shark is 1 cm longer than the base. The area is 21 cm^2. Find the height and base.
.
Let x = base
then
x+1 = height
.
area for any triangle is:
(1/2)bh = a
(1/2)x(x+1)= 21
multiplying both sides by 2:
x(x+1)= 42
x^2+x = 42
x^2+x-42=0
(x+7)(x-6) = 0
x = {-7,6}
we can throw out the negative solution (extraneous) leaving:
x = 6 cm (base)
.
height:
x+1 = 6+1 = 7 cm